# Modeling the formation of N2 and CH4 frost on Pluto’s slopes

### v1

@@ 2.3.1 Description of the parameterization for the Pluto VTM
The slope parameterization in the Mars Planetary Climate Model proposed by Lange et al. (2023) relies on the assumption that the annual mean solar irradiance of a terrain with a slope angle $\theta$ varies linearly with the cosine of the slope azimuth $\psi$. Hence, they concluded given slope ($\theta$, $\psi$) can be thermally, on average, represented by a slope with a slope angle of $\mid\mu\mid$ that is either North-facing if $\mu\penalty 10000\ >\penalty 10000\ 0$, or South-facing if $\mu\penalty 10000\ <\penalty 10000\ 0$, where $\mu$ is the projected slope on the meridional direction:
$\mu=\theta\cos(\psi)$ (1)
@@
However, this conclusion must be qualified for East-West oriented slopes on equatorial regions on Mars. The linear correlation observed between annual-mean solar irradiance on a slope and the projected slope $\mu$ at high latitudes is primarily due to the preferential azimuthal position of the Sun throughout the Martian year. At lower latitudes—i.e., within the tropics—Mars’ obliquity (25.2°) causes the solar declination to alternate between the Northern and Southern celestial hemispheres. This seasonal symmetry disrupts the dominance of a single solar azimuth and reduces the strength of the correlation between annual irradiance and $\mu$. Solar irradiance calculations for 30° slopes at various latitudes show that the correlation with $\mu$ holds only above approximately 10° latitude, with the values of the coefficient of determination, R2, dropping below 0.9 closer to the equator. Hence, the approximation proposed by Lange et al. (2023)—that the annual thermal behavior of a given slope ($\theta,\psi$) can be represented by its projected slope $\mu$—remains valid for present-day Mars at latitudes outside the equatorial band.
@@
The latter conclusion can not be directly applied to Pluto. Given its high obliquity -present-day value of 119.6°-, and thereby the high latitudes of the tropics ($\sim$ 60°), the correlation between the yearly-solar irradiance and slope azimuth is not straightforward on most of Pluto’s surface. When computing the solar irradiance for several steep slopes (Figure 1), we observe that the correlation is only clear for high latitudes ($\geq$ 70°), and disappears then rapidly at lower latitudes. Below this threshold, slopes with a significant east–west component cannot be thermally approximated by their projected slope $\mu$.
@@
Figure 1: Annual mean solar irradiance (at the surface) for several slope angles $\theta$ and azimuths $\psi$ ($\psi$ = 0° corresponding to a slope oriented Northward) at a latitude of 70°N (black), 30°N (green), 15°N (red), and 0°(blue). The correlation between the annual mean solar irradiance and the projected slope $\mu$ for each latitude is given by the coefficient of determination R2.
@@
Based on these considerations, our parameterization is restricted to the evolution of CH4 and N2 frost on north–south-oriented slopes, as slopes with an east–west component are likely not accurately represented. To implement this, each grid cell of the Pluto VTM is divided into a set of sub-grid units representing either North-facing slopes, South-facing slopes, or flat terrain (Figure 2). This simplification is justified by 1) the preferential accumulation of CH4 frost on pole-facing slopes at low latitudes (Earle et al., 2022, section 3.1), and 2) the fact that explicitly resolving all slope orientations would require introducing a larger number of sub-grid slope categories, leading to a prohibitive increase in computational cost.
@@
Figure 2: Schema of the sub-grid scale slope parameterization. Each coarse mesh of the global Pluto Volatile Transport Model/Pluto Planetary Climate Model (left) is decomposed into North-South facing sub-grid slopes (defined by characteristic slopes $\mu_{i}$) (right) or flat terrain. Each of these sub-grid surfaces is associated with its cover fraction $\delta_{i}$, which represents the percentage of the mesh grid occupied by a given slope $\mu_{i}$ within the observed local topography (middle). These sub-grid terrains have their own microclimate (insolation, surface and sub-surface temperatures, condensation/sublimation of frosts, etc.), and the interactions between the atmosphere and surface are made through averaged values over the mesh. On the right, bluish surfaces mean colder slopes than flat surfaces (and warmer for reddish surfaces).
@@
We use seven sub-grid slopes, three North-facing, three South-facing, and one flat surface, with $\mid{\mu_{1}}\mid\penalty 10000\ =\penalty 10000\ \mid{\mu_{7}}\mid$ = 30°, $\mid{\mu_{2}}\mid\penalty 10000\ =\penalty 10000\ \mid{\mu_{6}}\mid$ = 20°, $\mid{\mu_{3}}\mid\penalty 10000\ =\penalty 10000\ \mid{\mu_{5}}\mid$ = 10° and $\mu_{4}$ = 0°. The number of seven sub-grid slopes is chosen from our Mars’experience (Lange et al., 2023), which showed that such values allow a correct representation of the distribution of slopes within a mesh, without increasing the computational time significantly. Within each grid cell, every sub-grid slope is assigned a fractional coverage $\delta_{i}$, representing the percentage of the cell’s area occupied by slopes whose projected inclination lies within $\mu_{i}\pm 5$°. The cover fraction associated with a given projected slope angle $\mu_{i}$ is computed as follows. At each point of the Pluto DEM, we calculate the local slope and azimuth, and derive the corresponding $\mu$ value using Eq.1. When the slope is predominantly oriented East–West—defined as an azimuth $\psi$ within $\pm$45° of either 90° or 270°—we assume it behaves similarly to a flat surface and assign $\mu=0$°. The resulting global distribution of $\mu$ is shown in Figure3a. From this distribution, we define the cover fraction $\delta_{i}$ as the proportion of terrain with $\mu$ values within $\mu_{i}\pm 5$°, normalized by the total number of samples (Figure 3b). This distribution is then applied uniformly to all grid cells in the VTM, regardless of geographic location. We adopt this approach to avoid introducing artificial biases that could arise from the incomplete coverage of the DEM—particularly in the Southern hemisphere, where no topographic data are available. Attempting to extrapolate or impose sub-grid slope distributions in these poorly constrained regions could lead to spurious volatile fluxes in the model, for instance, by artificially channeling volatiles from smooth to rougher mesh cells. While it is likely that some regions of Pluto have indeed more slopes than others, using a single, globally averaged distribution ensures that such modeling artifacts are avoided and that simulations remain internally consistent despite the limitations of the topographic dataset. Moreover, because the VTM does not resolve atmospheric circulation and instead applies a uniform relaxation toward equilibrium, volatile transport remains simplified . This simplification means that potential transfers from rough to smooth terrains are not explicitly captured, which could otherwise bias flux estimates. We therefore retain a single, global slope distribution, and our results should be interpreted as showing where N2 or CH4 frost can form on a given slope rather than providing an accurate quantification of lateral volatile fluxes or total mass budgets for the frosts/ice.
@@
Figure 3: A) Normalized distribution of projected slopes $\mu$ — where $\mu$ = 0° corresponds to flat and East-West-facing slopes — on Pluto (solid grey histogram), Mars (red, from Lange et al., 2023), and Triton (derived from the fractional data of Schenk et al., 2021). B) Cover fraction $\delta_{i}$ derived from panel A), obtained by rebinning the histogram into the $\mu_{i}$ bins defined in the main text, each with a width of $\pm$ 5°.
@@
The slope surface temperature is computed based on its energy budget given by:
@@
$\begin{split}\rho c_{s}\frac{\partial T_{\rm{surf,slope}}}{\partial t}&=(1-A_{\rm{slope}})F_{\rm{rad,sw}}(t)+\epsilon_{\rm{slope}}F_{\rm{rad,lw}}(t)+F_{\rm{ground}}(t)\\ &\quad+\sum_{i}L_{i}\frac{\partial m_{i}}{\partial t}(t)-\epsilon_{\rm{slope}}\sigma T_{\rm{surf,slope}}^{4}(t)\end{split}$ (2)
@@
Assuming that the atmosphere is transparent at all wavelengths, and allowing only reflection of the solar radiation between the slope and the surrounding plains, the total solar irradiance on a given sub-grid slope $F_{\rm{rad,sw}}$ is given by:
$F_{\rm{rad,sw}}(t)=\frac{S_{0}}{R(t)^{2}}\left(\sin(\theta_{\rm{slope}}(t))+A_{\rm{flat}}(t)\sigma_{s}\sin(\theta_{\rm{flat}}(t))\right)$ (3)
where $S_{0}$ = 1361 W m-2 is the solar constant, $R$ is the distance between the Sun and Pluto (in astronomical units), $\theta_{\rm{slope}}$ is the angle of the Sun above the sloped surface (when $\sin(\theta_{\rm{slope}}(t))<0$, i.e., when the Sun if below the horizon, or the slope is in a self-shadow, then we set $\sin(\theta_{\rm{slope}}(t))$ to 0); and $\theta_{\rm{flat}}$ is the same angle but for the flat surface, $A_{\rm{flat}}$ is the albedo of the flat surface, and $\sigma_{s}$ (unitless) is the sky-view factor, which quantifies the proportion of the sky in the half hemisphere seen by the slope that is not obstructed by the surrounding terrain, computed with:
$\sigma_{s}=\frac{1+\cos(\mid\mu\mid)}{2}$ (4)
@@
Assuming again that the atmosphere is transparent, $F_{\rm{rad,lw}}(t)$ is then just computed with:
$F_{\rm{rad,lw}}(t)=(1-\sigma_{s})\epsilon_{\rm{flat}}(t)\sigma T_{\rm{surf,flat}}(t)^{4}$ (5)
@@
The computation of the ground flux $F_{\rm{ground}}$ is based on the 1D soil model originally described by Hourdin et al. (1993). To adequately resolve Pluto’s diurnal/seasonal scale lengths, the subsurface is divided into 24 discrete layers, with a geometrically stretched distribution, with the shallowest layer at 1.4 10-4 m and the deepest layer depth at nearly 1000 m.
@@
The amount of N2 frost condensing/subliming on a slope $\frac{\partial m_{\rm{N}_{2}}}{\partial t}$ (kg m-2) is derived from the amount of latent heat required to keep temperatures at the frost point if N2 ice is present, i.e
$\frac{\partial m_{\rm{N}_{2}}}{\partial t}(t)=\frac{\rho c_{s}}{\Delta tL_{\rm{N}_{2}}}\left(T_{\rm{cond}}-T_{0}\right)$ (6)
@@
where $\Delta t$ (s) is the timestep of the model, $L_{\rm{N}_{2}}$ = 2.5 105 J kg-1 is the latent heat of condensation for nitrogen, $T_{\rm{cond}}$ nitrogen condensation temperature derived from the thermodynamic relations of Fray & Schmitt (2009), and $T_{0}$ is the surface temperature on a sub-grid slope predicted by radiative and conductive balance $T_{0}$ (i.e., Eq. 2 solved without any condensation/sublimation of N2). This term is set to zero when no ice is present and $T_{0}>T_{\rm{cond}}$. If the amount of ice subliming is too high compared to what is available at the surface, then $\frac{\partial m_{\rm{N}_{2}}}{\partial t}(t)$ is set to the amount of ice available divided by the timestep.
The amount of CH4 frost condensing/subliming on a slope is given by:
$\frac{\partial m_{\rm{CH}_{4}}}{\partial t}(t)=\rho C_{d}U\left(q-q_{\rm{surf}}\right)$ (7)
@@
where $\rho$ (kg m-3) is the near-surface air density, $C_{d}$ (unitless) is a near-surface drag coefficient set to 0.06 (Forget et al., 2017), $U$ is the wind velocity near the surface and is set to 0.5 m s-1 (based on 3D GCM studies), $q$ (in kg kg-1) the CH4 atmospheric mass mixing ratio near the surface, and $q_{\rm{surf}}$ (in kg kg-1) is the saturation vapor pressure mass mixing ratio at the surface, computed using the thermodynamic Claudius–Clapeyron relation of Fray & Schmitt (2009) for CH4 with a latent heat for sublimation $L_{\rm{CH}_{4}}$ = 586.7 kJ kg-1. Again, if the latter equation predicts CH4 ice sublimation and no ice at the surface, this term is set to zero.
@@
Equation 2 is solved using an implicit scheme. However, applying such a scheme is non-trivial, as the surface temperature is intrinsically coupled to subsurface heat conduction—linking the surface radiative budget to the thermal diffusion within the ground (see for instance the mathematical derivations by Young, 2012). To address this, we adopt the strategy proposed by (Hourdin et al., 1995): the internal energy of the surface layer (and thus the surface temperature) is first updated at time $t$ using Equation 2. The temperatures of the subsurface layers are then computed by solving the heat conduction equation. Once these temperatures are known, the resulting soil heat flux is calculated and used as an input to Equation 2 at the next time step. This iterative procedure yields a stable and fully implicit scheme for our model.
@@
Once each sub-grid slope variables (e.g., slope temperature, albedo, etc.) are computed, grid parameters are determined by averaging all sub-grid parameter values weighted by their respective cover fractions. The grid box albedo and emissivity can be computed with :
$X=\sum_{i=1}^{7}X_{i}\delta_{i}$ (8)
@@
where $X$ refers to the grid parameter, $X_{i}$ the sub-grid parameter of slope $i$, $\delta_{i}$ the cover fraction.
@@
For the amount of volatile in the grid-box mesh, one must consider that the amount of ice on a sub-grid surface is expressed as a surface mass density (kg m-2). However, the area of the sloped sub-grid surface can be higher if it is sloped or flat, assuming a similar latitude/longitude boundary for the cell. Hence, to ensure mass conservation, the grid-box amount of volatile of a species $X$ in is computed with :
$X=\sum_{i=1}^{7}X_{i}\frac{\delta_{i}}{\cos(\mu_{i})}$ (9)
@@
Finally, the grid-box surface temperature is derived by averaging the Stefan–Boltzmann functions:
$\epsilon\sigma T_{\rm{surf,grid}}^{4}=\sum_{i=1}^{7}\epsilon_{i}\sigma T_{\rm{surf,i}}^{4}\delta_{i}$ (10)
@@
where $\epsilon$ (unitless) is the grid emissivity, $T_{\rm{surf,grid}}$ (K) the grid surface temperature, and quantities with subscript $i$ refer to the same quantities but for the sub-grid slope.
2.3.2 Modifications of the parameterization for the Pluto Planetary Climate Model

### v2

@@ 2.3.1 Description of the parameterization for the Pluto VTM
The slope parameterization in the Mars Planetary Climate Model proposed by Lange et al. (2023) relies on the assumption that the annual mean solar irradiance of a terrain with a slope angle $\theta$ varies linearly with the cosine of the slope azimuth $\psi$. Hence, they concluded given slope ($\theta$, $\psi$) can be thermally, on average, represented by a slope with a slope angle of $\mid\mu\mid$ that is either north-facing if $\mu\penalty 10000\ >\penalty 10000\ 0$, or south-facing if $\mu\penalty 10000\ <\penalty 10000\ 0$, where $\mu$ is the projected slope in the meridional direction,
$\mu=\theta\cos(\psi).$ (1)
@@
This conclusion must be qualified for east-west oriented slopes in equatorial regions on Mars, however. The linear correlation observed between annual-mean solar irradiance on a slope and the projected slope $\mu$ at high latitudes is primarily due to the preferential azimuthal position of the Sun throughout the Martian year. At lower latitudes (i.e., within the tropics), the obliquity of Mars (25.2°) causes the solar declination to alternate between the northern and southern celestial hemispheres. This seasonal symmetry disrupts the dominance of a single solar azimuth and reduces the strength of the correlation between annual irradiance and $\mu$. Solar irradiance calculations for 30° slopes at various latitudes show that the correlation with $\mu$ holds only above approximately 10° latitude, with the values of the determination coefficient, R2, dropping below 0.9 closer to the equator. Hence, the approximation proposed by Lange et al. (2023) that the annual thermal behavior of a given slope ($\theta,\psi$) can be represented by its projected slope $\mu$ remains valid for present-day Mars at latitudes outside the equatorial band.
@@
The latter conclusion cannot be directly applied to Pluto. Given its high obliquity -present-day value of 119.6°-, and thereby the high latitudes of the tropics ($\sim$ 60°), the correlation between the yearly-solar irradiance and slope azimuth is not straightforward on most of the Pluto surface. When computing the solar irradiance for several steep slopes (Fig. 1), we observe that the correlation is only clear for high latitudes ($\geq$ 70°), and disappears then rapidly at lower latitudes. Below this threshold, slopes with a significant east–west component cannot be thermally approximated by their projected slope $\mu$.
@@
Figure 1: Annual mean solar irradiance (at the surface) for several slope angles $\theta$ and azimuths $\psi$ ($\psi$ = 0° , corresponding to a slope oriented northward, at a latitude of 70°N (black), 30°N (green), 15°N (red), and 0° (blue). The correlation between the annual mean solar irradiance and the projected slope $\mu$ for each latitude is given by the determination coefficient R2.
@@
Based on these considerations, our parameterization is restricted to the evolution of CH4 and N2 frost on north–south-oriented slopes, as slopes with an east–west component are likely not accurately represented. To implement this, each grid cell of the Pluto VTM is divided into a set of subgrid units representing either north-facing slopes, south-facing slopes, or flat terrain (Fig. 2). This simplification is justified by 1) the preferential accumulation of CH4 frost on pole-facing slopes at low latitudes (Earle et al., 2022, Sect. 3.1), and 2) the fact that explicitly resolving all slope orientations would require introducing a larger number of subgrid-slope categories, leading to a prohibitive increase in computational cost.
@@
Figure 2: Schema of the subgrid-scale slope parameterization. Each coarse mesh of the global Pluto Volatile Transport Model/Pluto Planetary Climate Model (left) is decomposed into north-south facing subgrid slopes (defined by characteristic slopes $\mu_{i}$; right) or flat terrain. Each of these subgrid surfaces is associated with its cover fraction $\delta_{i}$, which represents the percentage of the mesh grid occupied by a given slope $\mu_{i}$ within the observed local topography (middle). These subgrid terrains have their own microclimate (insolation, surface, subsurface temperatures, condensation/sublimation of frosts, etc.), and the interactions between the atmosphere and surface are made through averaged values over the mesh. The bluish surfaces on the right show slopes that are colder than the flat surfaces (and warmer for reddish surfaces).
@@
We use seven subgrid slopes, three north-facing, three south-facing, and one flat surface, with $\mid{\mu_{1}}\mid\penalty 10000\ =\penalty 10000\ \mid{\mu_{7}}\mid$ = 30°, $\mid{\mu_{2}}\mid\penalty 10000\ =\penalty 10000\ \mid{\mu_{6}}\mid$ = 20°, $\mid{\mu_{3}}\mid\penalty 10000\ =\penalty 10000\ \mid{\mu_{5}}\mid$ = 10° and $\mu_{4}$ = 0°. The number of seven subgrid slopes is chosen from our Mars’experience (Lange et al., 2023), which showed that such values allow a correct representation of the distribution of slopes within a mesh, without increasing the computational time significantly. Within each grid cell, every subgrid slope is assigned a fractional coverage $\delta_{i}$, representing the percentage of the cell area occupied by slopes whose projected inclination lies within $\mu_{i}\pm 5$°. The cover fraction associated with a given projected slope angle $\mu_{i}$ is computed as follows. At each point of the Pluto DEM, we calculate the local slope and azimuth, and derive the corresponding $\mu$ value using Eq.1. When the slope is predominantly oriented east–west (defined as an azimuth $\psi$ within $\pm$45° of either 90° or 270°), we assume it behaves similarly to a flat surface and assign $\mu=0$°. The resulting global distribution of $\mu$ is shown in Fig.3a. From this distribution, we define the cover fraction $\delta_{i}$ as the proportion of terrain with $\mu$ values within $\mu_{i}\pm 5$°, normalized by the total number of samples (Fig. 3b). This distribution is then applied uniformly to all grid cells in the VTM, regardless of geographic location. We adopt this approach to avoid introducing artificial biases that could arise from the incomplete coverage of the DEM, particularly in the southern hemisphere, where no topographic data are available. Attempting to extrapolate or impose subgrid-slope distributions in these poorly constrained regions could lead to spurious volatile fluxes in the model, for instance, by artificially channeling volatiles from smooth to rougher mesh cells. While it is likely that some regions of Pluto have indeed more slopes than others, using a single, globally averaged distribution ensures that such modeling artifacts are avoided and that simulations remain internally consistent despite the limitations of the topographic dataset. Moreover, because the VTM does not resolve atmospheric circulation and instead applies a uniform relaxation toward equilibrium, volatile transport remains simplified . This simplification means that potential transfers from rough to smooth terrains are not explicitly captured, which could otherwise bias flux estimates. We therefore retain a single, global slope distribution, and our results should be interpreted as showing where N2 or CH4 frost can form on a given slope rather than providing an accurate quantification of lateral volatile fluxes or total mass budgets for the frosts/ice.
@@
Figure 3: (A): Normalized distribution of projected slopes $\mu$, where $\mu$ = 0° corresponds to flat and east-west-facing slopes, on Pluto (solid gray histogram), Mars (red, from Lange et al., 2023), and Triton (derived from the fractional data of Schenk et al., 2021). (B): Cover fraction $\delta_{i}$ derived from panel A), obtained by rebinning the histogram into the $\mu_{i}$ bins defined in the main text, each with a width of $\pm$ 5°.
@@
The slope surface temperature is computed based on its energy budget given by
@@
$\begin{split}\rho c_{s}\frac{\partial T_{\rm{surf,slope}}}{\partial t}&=(1-A_{\rm{slope}})F_{\rm{rad,sw}}(t)+\epsilon_{\rm{slope}}F_{\rm{rad,lw}}(t)+F_{\rm{ground}}(t)\\ &\quad+\sum_{i}L_{i}\frac{\partial m_{i}}{\partial t}(t)-\epsilon_{\rm{slope}}\sigma T_{\rm{surf,slope}}^{4}(t)\end{split},$ (2)
@@
Assuming that the atmosphere is transparent at all wavelengths, and allowing only reflection of the solar radiation between the slope and the surrounding plains, the total solar irradiance on a given subgrid slope $F_{\rm{rad,sw}}$ is given by
$F_{\rm{rad,sw}}(t)=\frac{S_{0}}{R(t)^{2}}\left(\sin(\theta_{\rm{slope}}(t))+A_{\rm{flat}}(t)\sigma_{s}\sin(\theta_{\rm{flat}}(t))\right),$ (3)
where $S_{0}$ = 1361 W m-2 is the solar constant, $R$ is the distance between the Sun and Pluto (in astronomical units), $\theta_{\rm{slope}}$ is the angle of the Sun above the sloped surface (when $\sin(\theta_{\rm{slope}}(t))<0$, i.e., when the Sun if below the horizon, or the slope is in a self-shadow, then we set $\sin(\theta_{\rm{slope}}(t))$ to 0); and $\theta_{\rm{flat}}$ is the same angle but for the flat surface, $A_{\rm{flat}}$ is the albedo of the flat surface, and $\sigma_{s}$ (unitless) is the sky-view factor, which quantifies the proportion of the sky in the half hemisphere seen by the slope that is not obstructed by the surrounding terrain, computed with
$\sigma_{s}=\frac{1+\cos(\mid\mu\mid)}{2},$ (4)
@@
Assuming again that the atmosphere is transparent, $F_{\rm{rad,lw}}(t)$ is then just computed with
$F_{\rm{rad,lw}}(t)=(1-\sigma_{s})\epsilon_{\rm{flat}}(t)\sigma T_{\rm{surf,flat}}(t)^{4}.$ (5)
@@
The computation of the ground flux $F_{\rm{ground}}$ is based on the 1D soil model originally described by Hourdin et al. (1993). To adequately resolve the diurnal/seasonal scale lengths on Pluto, the subsurface is divided into 24 discrete layers, with a geometrically stretched distribution, with the shallowest layer at 1.4 10-4 m and the deepest layer depth at nearly 1000 m.
@@
The amount of N2 frost condensing/subliming on a slope $\frac{\partial m_{\rm{N}_{2}}}{\partial t}$ (kg m-2) is derived from the amount of latent heat required to keep temperatures at the frost point if N2 ice is present, that is,
$\frac{\partial m_{\rm{N}_{2}}}{\partial t}(t)=\frac{\rho c_{s}}{\Delta tL_{\rm{N}_{2}}}\left(T_{\rm{cond}}-T_{0}\right),$ (6)
@@
where $\Delta t$ (s) is the timestep of the model, $L_{\rm{N}_{2}}$ = 2.5 105 J kg-1 is the latent heat of condensation for nitrogen, $T_{\rm{cond}}$ nitrogen condensation temperature derived from the thermodynamic relations of Fray and Schmitt (2009), and $T_{0}$ is the surface temperature on a subgrid slope predicted by radiative and conductive balance $T_{0}$ (i.e., Eq. 2 solved without any condensation/sublimation of N2). This term is set to zero when no ice is present and $T_{0}>T_{\rm{cond}}$. If the amount of ice subliming is too high compared to what is available at the surface, then $\frac{\partial m_{\rm{N}_{2}}}{\partial t}(t)$ is set to the amount of ice available divided by the timestep.
The amount of CH4 frost condensing/subliming on a slope is given by
$\frac{\partial m_{\rm{CH}_{4}}}{\partial t}(t)=\rho C_{d}U\left(q-q_{\rm{surf}}\right),$ (7)
@@
where $\rho$ (kg m-3) is the near-surface air density, $C_{d}$ (unitless) is a near-surface drag coefficient set to 0.06 (Forget et al., 2017), $U$ is the wind velocity near the surface and is set to 0.5 m s-1 (based on 3D GCM studies), $q$ (in kg kg-1) the CH4 atmospheric mass mixing ratio near the surface, and $q_{\rm{surf}}$ (in kg kg-1) is the saturation vapor pressure mass mixing ratio at the surface, computed using the thermodynamic Claudius–Clapeyron relation of Fray and Schmitt (2009) for CH4 with a latent heat for sublimation $L_{\rm{CH}_{4}}$ = 586.7 kJ kg-1. Again, if the latter equation predicts CH4 ice sublimation and no ice at the surface, this term is set to zero.
@@
Equation 2 is solved using an implicit scheme. It is nontrivial to apply this scheme, however, because the surface temperature is intrinsically coupled to subsurface heat conduction, which links the surface radiative budget to the thermal diffusion within the ground (see for instance the mathematical derivations by Young, 2012). To address this, we adopt the strategy proposed by (Hourdin et al., 1995): the internal energy of the surface layer (and thus the surface temperature) is first updated at time $t$ using Eq. 2. The temperatures of the subsurface layers are then computed by solving the heat conduction equation. Once these temperatures are known, the resulting soil heat flux is calculated and used as an input to Eq. 2 at the next time step. This iterative procedure yields a stable and fully implicit scheme for our model.
@@
Once each subgrid-slope variables (e.g., slope temperature, and albedo) are computed, grid parameters are determined by averaging all subgrid parameter values weighted by their respective cover fractions. The grid box albedo and emissivity can be computed with
$X=\sum_{i=1}^{7}X_{i}\delta_{i},$ (8)
@@
where $X$ refers to the grid parameter, $X_{i}$ the subgrid parameter of slope $i$, $\delta_{i}$ the cover fraction.
@@
For the amount of volatile in the grid-box mesh, one must consider that the amount of ice on a subgrid surface is expressed as a surface mass density (kg m-2). The area of the sloped subgrid surface can be higher if it is sloped or flat, however, when a similar latitude/longitude boundary is assumed for the cell. Hence, to ensure mass conservation, the grid-box amount of volatile of a species $X$ in is computed with
$X=\sum_{i=1}^{7}X_{i}\frac{\delta_{i}}{\cos(\mu_{i})}.$ (9)
@@
Finally, the grid-box surface temperature is derived by averaging the Stefan–Boltzmann functions
$\epsilon\sigma T_{\rm{surf,grid}}^{4}=\sum_{i=1}^{7}\epsilon_{i}\sigma T_{\rm{surf,i}}^{4}\delta_{i},$ (10)
@@
where $\epsilon$ (unitless) is the grid emissivity, $T_{\rm{surf,grid}}$ (K) the grid surface temperature, and quantities with subscript $i$ refer to the same quantities but for the subgrid slope.
2.3.2 Modifications of the parameterization for the planetary climate model for Pluto